OpenStudy (anonymous):

A sequence is defined by the formula f(n + 1) = f(n) – 3. If f(4) = 22, what is f(1)?

OpenStudy (anonymous):

d=f(n+1)-f(n)=-3 an=a1+(n-1)d 22=f(1)+(4-1)*(-3) ?

OpenStudy (danjs):

this seems alright to read, then you can finish your work much faster

OpenStudy (capnkookcx):

https://soundcloud.com/dann-gooas/record-0049

jaynator495 (jaynator495):

Welcome To OpenStudy! Here you will find great helpers and friends, a community of students that help students! We hope you enjoy the experience!